OpenStudy (anonymous):

[3.06] Given the function f(x) = 2(x + 4), find x if f(x) = 20. 6 48 14 8

OpenStudy (ash2326):

f(x)=2(x+4) f(x)=20 2(x+4)=20 Divide both sides by 2 \[2\frac{x+4}{2}=\frac{20}{2}\] \[x+4=10\] Can you find x now ? @swigglb

OpenStudy (anonymous):

thnxxx

OpenStudy (anonymous):

got it

OpenStudy (ash2326):

Now do the same thing for all parts

OpenStudy (anonymous):

got another question

OpenStudy (anonymous):

If U = {positive integers} and subset M = {positive odd integers}, what is M'?